How much of each solution, a 30% solution and a 50% solution, would be added together to produce 10 litres of a 45% solution?
Found 2 solutions by macston, ikleyn:
Answer by macston(5194)

.
T=amount of 30% solution; F=amount of 50% solution
.
T+F=10L
T=10L-F
.
0.30T+0.50F=0.45(10L)
0.30(10L-F)+0.50F=4.5L
3.0L-0.30F+0.50F=4.5L
0.20F=1.5L
F=7.5L
.
T=10L-F=10L-7.5L=2.5L
.
ANSWER: Add 2.5 liters of 30% solution and 7.5 liters of
50% solution to produce 10 liters of 45% solution.

Answer by ikleyn(52787)

.
See the lesson Solving typical word problems on mixtures for solutions in this site.

It explains how to solve mixture problems similar to this one.